# CuckooGraph: A Scalable and Space-Time Efficient Data Structure for Large-Scale Dynamic Graphs

### v1

@@ V-A Experimental Setup
Implementation: We implement CuckooGraph and the other competitor algorithms with C++ and build them with g++ 7.5.0 and -O3 option. The hash functions we use are 32-bit Bob Hash (obtained from the open-source website [33]) with random initial seeds. For CuckooGraph, we set $R=3$, the Denylist size to 128, as well as the ratio of the number of buckets in the two arrays of L/S-CHT is 2:1, and whether the basic or extended version of CuckooGraph is used depends on whether the dataset has repeated edges. For WBI, we set the length and width of its ceiling/adjacency matrix to 400 for better query performance.
Datasets: We use four real-world datasets and one synthetic datasets: 1) CAIDA; 2) NotreDame; 3) StackOverflow; 4) WikiTalk; 5) DenseGraph. The details are shown as follows.
1) CAIDA: The 1st dataset is streams of anonymized IP traces collected by CAIDA [34]. Each flow is identified by a five-tuple: source and destination IP addresses, source and destination ports, protocol. The source and destination IP addresses in the traces are used as the start and end points of the graph, respectively. The dataset contains 27121713 duplicate edges concerning 510839 nodes, with an edge density of $3.26\times 10^{-6}$ and an average node degree of 1.66.
2) NotreDame: The 2nd dataset is a web graph collected from University of Notre Dame [35]. Nodes represent web pages, and directed edges represent hyperlinks between them. The dataset contains 1497134 duplicate-free edges concerning 325729 nodes, with an edge density of $1.41\times 10^{-5}$ and an average node degree of 4.60.
3) StackOverflow: The 3rd dataset is a collection of interactions on the stack exchange website called Stack Overflow [36]. Nodes represent users and edges represent user interactions. The dataset has 63497050 duplicate edges and 2601977 nodes, with an edge density of $5.35\times 10^{-6}$ and an average node degree of 13.9.
4) WikiTalk: The 4th dataset is a collection of user communications obtained from English Wikipedia [37]. The nodes and edges refer to the same as above. It includes 24981163 duplicate edges and 2987535 nodes, with an edge density of $1.05\times 10^{-6}$ and an average node degree of 3.14.
5) DenseGraph: We synthesize this large dense graph dataset consisting of many simple directed graphs. It has 57593245 duplicate-free edges and 8000 nodes, with an edge density of $0.90$ and an average node degree of 7199.2.
Evaluation Metrics: We use the following metrics to compare the performance of our CuckooGraph with the SOTA competitors.
1) Average Insertion Time (AIT): AIT is defined as the average time required to insert an edge.
2) Average Query Time (AQT): AQT is defined as the average time required to perform a given query task.
3) Memory Usage (MU): MU is defined as the memory used to store a specified amount of edges.
Legend Note: We provide explanations of the abbreviations in the legends of some figures in Table III.
V-B Experiments on Parameter Settings
In this section, we measure the effects of some key parameters for CuckooGraph, namely, the number of cells per bucket in L/S-CHT d, the preset $LR$ threshold for expansion G, and the maximum number of loops in L/S-CHT T. This experiment evaluates the effects by: 1) We first batch inserting all edges in the CAIDA dataset into CuckooGraph and then batch querying them from CuckooGraph, and measure the average throughput separately; 2) We measure the memory usage by continuously inserting edges.
@@ V-B Experiments on Parameter Settings
Effects of $G$ (Figure 3(a))-3(c): Our experimental results show that the overall performance is best when the value of $G$ is 0.9. We find that the insertion and query throughput of CuckooGraph with $G$ of 0.8, 0.85, and 0.9 are very close to each other, and all are faster than the one at $G$ of 0.95. In addition, the larger $G$ is, the smaller the memory usage of CuckooGraph is. Thus, we set $G=0.8$ after the above considerations.
@@
V-C Experiments on Loading Rate
We conduct this loading rate experiment to verify that a single S-CHT can achieve a very high loading rate without insertion failure under our settings. For example, assuming that the first insertion failure occurs when the loading rate is 0.95, the probability of insertion failure when the loading rate is below 0.8 is extremely low (no Denylist is involved). Specifically, given the length of S-CHT, we use $v$ of the dataset to continuously insert S-CHT until the first insertion failure occurs, and output the loading rate of the S-CHT at this moment.
Figure 5: Evaluation on Loading Rate When First Insertion Fails.
Results & Analysis (Figure 5): On different datasets and S-CHTs of different lengths, the loading rate can almost always reach above 0.95 (commonly to reach above 0.98) before the first insertion failure occurs. It shows that our S/L-CHT with two bucket array length ratio of 2:1 and $d=8$ can achieve extremely high loading rate, and the probability of insertion failure occurring below our setting of $G=0.8$ is extremely low.
V-D Experiments on Expansion Time
CuckooGraph’s S-CHT faces expansion during the insertion process, and only the Transformation from the 2n/n/n form to the 4n/2n form is the most time-consuming. The above time-consuming mainly lies in the operation of merging three 2n/n/n tables into one 4n table. Therefore, we use this experiment to explore how time-consuming such a merging operation is. Specifically, we use the deduplicated $v$ to continuously insert into these 3 empty tables so that their loading rate reach 0.8, then merge them and calculate the time spent on the merging operation.
Figure 6: Evaluation on the Time Required for Expansion Operations.
Results & Analysis (Figure 6): The merging time has an approximately linear relationship with the S-CHT length. Although one merging may seem slightly longer, such time consumption is acceptable. Since the frequency of merging is very low, considering that merging only occurs once every many insertions, the AIT can still be stably maintained at a low value. Our Theorem 2 has shown that the AIT does not increase as the amount of data increases, but remains constant.
V-E Experiments on Edge Queries
@@
Methodology: 1) We insert a certain number of edges from the dataset into CuckooGraph or WBI and get the total insertion time, calculate it divided by the number of insertions to get the AIT. 2) We query the inserted edges sequentially and get the total query time and calculate it divided by the number of queries to get the AQT. 3) We first deduplicate the dataset with duplicate edges and obtain non-duplicated edges, and then insert them into CuckooGraph or WBI in sequence. After each insertion, the memory usage of CuckooGraph or WBI at that moment is output in sequence.
AIT (Figure 7(a)-7(d)): We find that the AIT of the NT and T states of CuckooGraph is around 24.0$\times$ and 21.6$\times$ lower than that of WBI on average, respectively.
AQT (Figure 8(a)-8(d)): We find that the AQT of the NT and T states of CuckooGraph is around 51.5$\times$ and 44.6$\times$ lower than that of WBI on average, respectively.
MU (Figure 9(a)-9(d)): We find that the MU of CuckooGraph at the three loading rates (0.8, 0.85, 0.9) is around 2.34$\times$, 2.46$\times$, and 2.56$\times$ lower than that of WBI on average, respectively.
Analysis: The time complexity of CuckooGraph insertion and query is $O(1)$ without considering the time-consuming changes of the hash function modulo operations. The maximum number of L-CHT and S-CHT per cell is 3, and the lengths of L-DL and S-DL are also limited. Assuming that accessing a bucket or a Denylist requires one memory access, the worst case scenario when querying an edge is no more than 14 memory accesses, so there is an upper bound for the number of memory accesses in CuckooGraph. We can also demonstrate that the AIT in the worst case is no more than 3 times the above insertion time for a single insertion that does not involve Transformation. Conversely, since the adjacency matrix size of WBI is fixed, its insertion and query time increase as the number of different inserted edges increases, and the fixed adjacency matrix makes it poorly adaptable to graph scale changes. In terms of space usage, WBI suffers from several important drawbacks. First, WBI has to store a large number of pointers. CuckooGraph does not need to store a large number of additional pointers, and it has few structures that are not directly used for storage. Its space waste comes mainly from L/S-CHTs not being filled, but this is compensated for by the fact that its upper limit on the loading rate can be very high. Second, in WBI, assuming that a $u$ corresponds to $k$ $v$, the $u$ has to be stored $k$ times in the $k$ linked list “blocks” corresponding to those $k$ different edges. When $k$ is large, such repeated storage imposes a large space overhead. Conversely, in CuckooGraph, unless $u$ has a corresponding $v$ that is kicked to S-DL, the $u$ need only be stored once and there is no duplicate storage. Third, WBI requires additional storage of additional data structures such as an adjacency matrix and counters. If the adjacency matrix is set too large in advance due to unpredictable data volume, it will cause huge waste of space. Conversely, CuckooGraph can flexibly handle any data volume.
@@
Figure 7: CuckooGraph v.s. WBI, Edge Query Performance on Average Insertion Time.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 8: CuckooGraph v.s. WBI, Edge Query Performance on Average Query Time.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 9: CuckooGraph v.s. WBI, Edge Query Performance on Memory Usage.
V-F Experiments on 1-Hop Successor and Precursor Queries
V-F1 1-Hop Successor Queries
Methodology: This type of query refers to giving a $u$ and returning all its $v$ and their number. For CuckooGraph, it means finding the cell ($l\leq 2R$) or all S-CHT ($l>2R$) corresponding to the $u$, and returning all $v$ among them. In addition, S-DL must also be traversed and queried in case there is a $v$ corresponding to the $u$. For WBI, we first find all the rows corresponding to the $u$ on the adjacency matrix, traverse all the edges hanging under these rows to find all the edges corresponding to the $u$ and return their corresponding $v$. We evaluate this by inserting a certain number of non-duplicate edges, after which we randomly pick 5000 of the inserted $u$ for successor queries and calculate the AQT.
AQT (Figure 10(a)-10(d)): We find that the AQT of the NT and T states of CuckooGraph is around 13500.8$\times$ and 9994.7$\times$ lower than that of WBI on average, respectively.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 10: CuckooGraph v.s. WBI, 1-Hop Successor Query Performance on Average Query Time.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 11: CuckooGraph v.s. WBI, 1-Hop Precursor Query Performance on Average Query Time.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 12: CuckooGraph v.s. WBI, Finding Triangles Performance on Average Query Time.
(a) CAIDA
(b) NotreDame
(c) StackOverflow
(d) WikiTalk
Figure 13: CuckooGraph v.s. WBI, Single Source Shorted Path Performance on Execution Time.
(a) AIT
(b) AQT
(c) MU
Figure 14: CuckooGraph v.s. WBI, Edge Query Performance on DenseGraph Dataset.
V-F2 1-Hop Precursor Queries
Methodology: This type of query is the opposite of the above-mentioned successor query, $i.e.$, given a $v$, all its $u$ and their number are returned. Note that for $\left\langle u,v\right\rangle$, we need to insert it into CuckooGraph according to $\left\langle v,u\right\rangle$, $i.e.$, insert it in reverse. For WBI, we can insert normally like the above successor, but when querying the predecessor, we need to change “traverse all rows corresponding to $u$” to “traverse all columns corresponding to $v$”.
AQT (Figure 11(a)-11(d)): We find that the AQT of the NT and T states of CuckooGraph is around 12839.8$\times$ and 9210.6$\times$ lower than that of WBI on average, respectively.
V-F3 Analysis
First, successor queries of WBI need to traverse a huge number of irrelevant other edges, resulting in a lot of waste of overhead. On the contrary, except that CuckooGraph may traverse a small number of edges of other $u$ in S-DL, the rest of the $v$ traversed are all based on the target $u$. In short, the amount it traverses is much smaller than that of WBI. Second, the linked list “blocks” used by WBI to store edges are connected by pointers. Therefore, the addresses of different blocks may be far apart, and the spatial locality is very poor, so that traversing WBI may require a huge amount of memory accesses. As for CuckooGraph, $v$ basically exists in a small number of S-CHTs. Therefore, different $v$ can be arranged much more continuously in address with good spatial locality, which can greatly reduce the number of memory accesses. The analysis of the precursor queries is similar to that of the successor queries and will not be described in detail.
V-G Experiments on Finding Triangles
Methodology: Finding triangles means that given a node, return the number of triangles containing the node in the graph. For this query, WBI first finds all the precursors and all successors of the node, and then enumerates all possible edges $\left\langle successor,precursor\right\rangle$ composed of these successors and precursors to perform edge queries. The number of successful queries is the result of this task. CuckooGraph first finds all the 2-hop successors of the node through successor queries, and then enumerates all the possible edges $\left\langle 2-hop\ successor,node\right\rangle$ composed of the node’s 2-hop successors and the node itself to perform edge queries. The evaluation method is also similar to Section V-F.
AQT (Figure 12(a)-12(d)): We find that the AQT of the NT and T states of CuckooGraph is around 136.5$\times$ and 75.2$\times$ lower than that of WBI on average, respectively.
Analysis: The reason is that CuckooGraph’s successor, precursor, and edge queries are all much faster than WBI’s. Because finding triangles is accomplished based on the successor, precursor and edge queries.
V-H Experiments on SSSP
Methodology: The details of the Single-Source Shortest Paths (SSSP) task are as follows. We first insert all the non-repeating edges of the dataset, and then select a number of nodes with the largest total degree ($i.e.$, the sum of the outgoing and incoming degrees) to extract the subgraph. After that, we take the 10 nodes with the largest total degree among these nodes, and execute Dijkstra algorithm [38] 10 times with these 10 nodes as the source. Each time Dijkstra algorithm is used to find the shortest distance from the source to the node on the subgraph for each node in the subgraph. We evaluate the execution time of these 10 tests. In Dijkstra algorithm, whenever an edge query is needed, an edge query of CuckooGraph or WBI is executed. Note that in this experiment, we default to a weight of 1 for all edges.
@@ V-H Experiments on SSSP
Execution Time (Figure 13(a)-13(d)): We find that the Execution Time of the NT and T states of CuckooGraph is around 132.8$\times$ and 103.7$\times$ lower than that of WBI on average, respectively.
Analysis: The reason why CuckooGraph is faster than WBI in this experiment is because CuckooGraph’s edge query is faster than WBI, especially when the amount of data is large.
V-I Other Evaluations
V-I1 From Real-World Graphs to Dense Graphs
The real-world datasets used in the experiments before this section are all sparse graphs. In order to verify that CuckooGraph is also effective on dense graphs, we synthesize the DenseGraph dataset and use AIT, AQT, and MU as the evaluation metrics.
Results & Analysis (Figure 14(a)-14(c)): 1) The AIT/AQT of the NT and T states of CuckooGraph is around 26.2×26.2$\times$26.2 ×/71.8×71.8$\times$71.8 × and 25.4×25.4$\times$25.4 ×/60.8×60.8$\times$60.8 × lower than that of WBI on average, respectively; 2) The MU of CuckooGraph at the three loading rates (0.8, 0.85, 0.9) is around 4.59$\times$ lower than that of WBI on average, respectively.
V-I2 A Discussion with Graph Stream Summarization
In recent years, the research community has proposed a series of graph stream summarization schemes [39, 40, 41, 42, 43, 44, 45] to approximately store graph streams in sublinear space, which is a different problem from the one concentrated in this paper. Next, we kindly compare and discuss three graph stream summarization schemes (Auxo [45], GSS [41], and TCM [40]) with the extended version of CuckooGraph (W-ours) on weighted edge queries in terms of AIT and AQT on CAIDA.
Results & Analysis (Figure 15(a)-15(b)): We find that both AIT and AQT of CuckooGraph are always faster than Auxo/GSS/TCM on weighted edge queries. However, these experiments are not comparisons between similar solutions, but are intended as friendly comparisons for discussion purposes only. It is clear that CuckooGraph/WBI records complete graph information, while Auxo/GSS/TCM approximately records graph streams, so they each have their own advantages and disadvantages and are designed for different scenarios and requirements.
(a) AIT
(b) AQT
Figure 15: Comparative experiments on the weighted edge query task.

### v2

@@ V-A Experimental Setup
Implementation: We implement CuckooGraph and the other competitors with C++ and build them with g++ 7.5.0 and -O3 option. The hash functions we use are 32-bit Bob Hash (obtained from the open-source website [45]) with random initial seeds. For CuckooGraph, we set $R=3$, as well as the ratio of the number of buckets in the two arrays of L/S-CHT is 2:1, and whether the basic or extended version of CuckooGraph is used depends on whether the dataset has repeated edges.
Competitors: Since there are many related works on dynamic graph storage, we rigorously select some of the SOTA ones from recent years for experimental comparison: LiveGraph [30], Sortledton [34], Wind-Bell Index (WBI) [35], and Spruce [36].
Datasets: We use various graph datasets to comprehensively evaluate the performance of CuckooGraph and its competitors, and the details are shown in Table III. 1) The CAIDA dataset [46] is streams of anonymized IP traces collected by CAIDA. Each flow is identified by a five-tuple: source and destination IP addresses, source and destination ports, protocol. The source and destination IP addresses in the traces are used as the start and end nodes of the graphs, respectively. 2) The NotreDame dataset [47] is a web graph collected from University of Notre Dame. Nodes represent web pages, and directed edges represent hyperlinks between them. 3) The StackOverflow dataset [48] is a collection of interactions on the stack exchange website called Stack Overflow. Nodes represent users and edges represent user interactions. 4) The WikiTalk dataset [49] is a collection of user communications obtained from English Wikipedia [49], and the nodes and edges refer to the same as above. 5) The Weibo dataset [50] is captured from Sina Weibo Open Platform APIs, and the definitions of nodes and edges are similar to those of StackOverflow. 6) We synthesize the DenseGraph dataset. 7) We synthesize the SparseGraph dataset.
Metrics: We use the following key metrics.
  • Throughput: It is defined as Million Operations Per Second (Mops). We use Throughput to evaluate the average insertion, query, and deletion speed.
  • Memory Usage: It is defined as the memory used to store a specified amount of edges.
  • Running Time: It is defined as the time spent performing the specified graph analytics tasks.
@@
V-B Experiments on Parameter Settings
In this subsection, we measure the effects of some key parameters for CuckooGraph, namely, the number of cells per bucket in L/S-CHT d, the preset $LR$ threshold for expansion G, and the maximum number of loops in L/S-CHT T. This experiment evaluates the effects by: 1) We first batch inserting edges in the CAIDA dataset into CuckooGraph and then batch querying them from CuckooGraph, and measure the average throughput separately; 2) We measure the memory usage by continuously inserting edges.
@@ V-B Experiments on Parameter Settings
Effects of $G$ (Figure 3(a))-3(c): Our experimental results show that the overall performance is best when the value of $G$ is 0.9. We find that the insertion and query throughput of CuckooGraph with $G$ of 0.8, 0.85, and 0.9 are very close to each other, and all are faster than the one at $G$ of 0.95. In addition, the larger $G$ is, the smaller the memory usage of CuckooGraph is. Thus, we set $G=0.9$ after the above considerations.
@@
V-C Ablation Experiments
In this subsection, we conduct ablation experiments to evaluate the individual effects of Denylist (DL) optimization on CuckooGraph performance to verify its effectiveness. Our methodology is that every time an insertion failure occurs, we expand the size of CuckooGraph to 1.5×1.5$\times$1.5 × its original size. We use the CAIDA dataset and evaluate the effects in terms of insertion and query throughput as well as memory usage.
(a) Insertion Throughput
(b) Query Throughput
(c) Memory Usage
Figure 5: Ablation experiments: CuckooGraph with and without DL optimization.
Effects of DL (Figure 5): The experimental results show that DL indeed further speeds up insertion and querying with almost no additional memory overhead. We find that the insertion and query throughput of CuckooGraph with DL optimization is $1\times1$ and 1.12×1.12$\times$1.12 × faster than that of CuckooGraph without DL optimization, respectively. Also, the memory usage of CuckooGraph with DL optimization is only about 4KB more than that of CuckooGraph without DL optimization when all items are inserted.
V-D Experiments on Throughput and Memory Usage
In this subsection, we evaluate the performance of CuckooGraph and its competitors in terms of insertion, query, and deletion throughput and memory usage on various graph datasets.
@@ V-D Experiments on Throughput and Memory Usage
Methodology: 1) We insert all edges from the graph dataset into an empty graph structure, and calculate the average insertion throughput; 2) We then query all edges from the graph structure and calculate the average query throughput. 3) We delete edges one by one and calculate the throughput of the process after deletions. 4) We first de-duplicate the datasets to obtain non-duplicated edges, and then insert them into each scheme one by one. After each insertion, the physical memory overhead at that moment is output.
Insertion throughput (Figure 6): The results show that, on the seven datasets, the insertion throughput of CuckooGraph is 72.17×72.17$\times$72.17 ×, 32.66×32.$6$, 8.60×8.60$\times$8.60 ×, and 253.32×253.32$\times$253.32 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Figure 6: Insertion throughput on different datasets.
Figure 7: Query throughput on different datasets.
Figure 8: Deletion throughput on different datasets.
Query throughput (Figure 7): The results show that, on the seven datasets, the query throughput of CuckooGraph is 14.69×14.69$\times$14.69 ×, 133.62×133.62$\times$133.62 ×, 5.34×5.34$\times$5.34 ×, and 287.48×287.48$\times$287.48 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Deletion throughput (Figure 8): The results show that, on the seven datasets, the deletion throughput of CuckooGraph is 85.47×85.47$\times$85.47 ×, 3.63×3.63$\times$3.63 ×, 5.01×5.01$\times$5.01 ×, and 65.55×65.$5$ faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: 1) Thanks to the novel data structure of CuckooGraph, when inserting or querying an edge, even in the worst case, only 6 buckets in L-CHT and S-CHT, as well as two Denylists, are accessed. Since the size of the bucket and Denylist is fixed, the upper limit on the number of memory accesses is also fixed and small. Therefore, no matter how the incoming dataset changes, CuckooGraph can achieve fast insertion and query. In contrast, other competitors are designed based on the adjacency list or its variants, so an edge insertion/query operation often requires multiple memory accesses and cannot adapt well to changes in the amount and characteristics of the dataset. 2) For deletions, other schemes simply delete the target when it is found, while CuckooGraph may involve additional contraction operations.
@@
(e) Weibo
(f) DenseGraph
(g) SparseGraph
Figure 9: Memory usage on different datasets.
Memory Usage (Figure 9(a)-9(g)): The results show that, on the seven datasets, the memory usage of CuckooGraph when all item insertions are completed is 5.92×5.92$\times$5.92 ×, 1.47×1.47$\times$1.47 ×, 4.89×4.89$\times$4.89 ×, and 2.34×2.34$\times$2.34 × less than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: CuckooGraph is a customized design based on CHT, so it does not need to store a large number of pointers like the schemes based on adjacency lists, which significantly reduces space overhead. In addition, since CHT has a high loading rate, CuckooGraph achieves a high loading rate and minimizes space waste.
V-E Experiments on Graph Analytics Tasks
In this subsection, we evaluate the performance of CuckooGraph and its competitors in terms of running time on the graph datasets in Table III through the following typical graph analytics tasks: Breadth-First Search (BFS), Single-Source Shortest Paths (SSSP), Triangle Counting (TC), Connected Components (CC), PageRank (PR), Betweenness Centrality (BC), and Local Clustering Coefficient (LCC). Note that some competitors did not complete the experiments within the given time, so their results are not shown in the provided figures.
V-E1 Breadth-First Search
Methodology: We first insert all the edges of the entire dataset. Then, we select a specific number of nodes with the largest total degree ($i.e.$, the sum of out-degree and in-degree, the same below), and perform a BFS on these nodes, returning each node and the number of nodes obtained in the order of BFS traversal. Finally, we calculate the average time taken for these BFS tasks.
@@ V-E1 Breadth-First Search
Results (Figure 10): We find that, on the seven datasets (two for WBI), the running time of CuckooGraph on BFS is 2.34×2.34$\times$2.34 ×, 0.73×0.73$\times$0.73 ×, 19.83×19.83$\times$19.83 ×, and 504.81×504.81$\times$504.81 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: The most frequently used function of each scheme in this task is its successor query function. The structure of CuckooGraph is based on hash tables, so it has good spatial locality. Therefore, during the process of querying and traversing the hash tables, the algorithm can achieve excellent spatial locality, which greatly increases the cache hit rate. Most other adjacency list-based schemes need to store data in different memory addresses and then use pointers to link them. When querying for successors, the time and space locality is poor and the cache hit rate is low, requiring frequent memory access, which reduces query efficiency. It is worth noting that WBI not only has the above shortcomings, but also needs to access many other redundant edges when querying successors, so it performs the worst. The advantage of Spruce may be that its end nodes for finding neighbors can be approximately regarded as being stored more continuously than the other 3.
Figure 10: Running time of BFS on different datasets.
Figure 11: Running time of SSSP on different datasets.
Figure 12: Running time of TC on different datasets.
Figure 13: Running time of CC on different datasets.
V-E2 Single-Source Shortest Paths
Methodology: We first insert all the edges of the entire dataset. Then, we select a specific number of nodes with the largest total degree to extract subgraphs, and select the 10 nodes with the largest total degree among these nodes. Note that this refers to the 10 nodes with the largest total degree on the original graphs, not on the subgraphs. After that, we use these 10 nodes as sources to perform Dijkstra algorithm [51] 10 times and calculate the average time.
Results (Figure 11): We find that, on the seven datasets (six for Spruce & WBI), the running time of CuckooGraph on SSSP is 43.64×43.64$\times$43.64 ×, 168.45×168.45$\times$168.45 ×, 1.62×1.62$\times$1.62 ×, and 278.0×278.0$\times$278.0 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: The most frequently used function of each scheme in this task is edge query function. As described in the analysis in §$\S$ V-D, CuckooGraph has a huge advantage over other adjacency list-based schemes in edge query, so CuckooGraph achieves the best performance in the SSSP task.
V-E3 Triangle Counting
[…]
Results (Figure 12): We find that, on the seven datasets (five for WBI), the running time of CuckooGraph on TC is 8.23×8.23$\times$8.23 ×, 21.33×21.$3$, 1.86×1.86$\times$1.86 ×, and 3015.11×3015.$1$ faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: The edge query and successor query functions are the most frequently used functions in each scheme in this task. As analyzed in §$\S$ V-E1 and §$\S$ V-E2, CuckooGraph can achieve good performance in these two functions, so it also performs very well in this task.
Figure 14: Running time of PR on different datasets.
Figure 15: Running time of BC on different datasets.
Figure 16: Running time of LCC on different datasets.
V-E4 Connected Components
Methodology: All edges for the entire dataset are inserted. We first select a specific number of nodes with the largest total degree to extract subgraphs, and then insert the subgraphs into each scheme. Note that the above steps also apply to the last 3 tasks. After that, we run the Tarjan algorithm [52] on the subgraphs using each scheme and return the connected components and their number.
Results (Figure 13): We find that, on the seven datasets, the running time of CuckooGraph on CC is 2.11×2.$1$, 1.07×1.07$\times$1.07 ×, 9.91×9.91$\times$9.91 ×, and 39.7×39.7$\times$39.7 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: The most frequently used function of each scheme in this task is its successor query function. As analyzed in §$\S$ V-E1, CuckooGraph performs well in this function, so it has an advantage over other schemes in this task.
V-E5 PageRank
Methodology: The initial steps are the same as those in §$\S$ V-E4. Then, we use the successor query function of each scheme to assist in constructing the matrix required to solve the PageRank (PR), and iterate 100 times on the matrix to find the PR of each node on the subgraphs.
Results (Figure 14): We find that, on the seven datasets, the running time of CuckooGraph on PR is 2.16×2.16$\times$2.16 ×, 1.03×1.03$\times$1.03 ×, 2.62×2.62$\times$2.62 ×, and 2.87×2.87$\times$2.87 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: Each scheme in this task frequently uses the successor query function to construct the matrix required to calculate PR. As analyzed in §$\S$ V-E1, CuckooGraph performs well in successor query, so it also shows advantages over other schemes in this task.
V-E6 Betweenness Centrality
Methodology: The initial steps are the same as those in §$\S$ V-E4. Then, we run the Brandes algorithm [53] on the subgraphs using each scheme.
Results (Figure 15): We find that, on the seven datasets, the running time of CuckooGraph on BC is 3.15×3.15$\times$3.15 ×, 16.17×16.17$\times$16.17 ×, 7.33×7.$3$, and 5.23×5.23$\times$5.23 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: Similar to the analysis in §$\S$ V-E4.
V-E7 Local Clustering Coefficient
Methodology: The initial steps are the same as those in §$\S$ V-E4. Then, we pre-compute all neighbors of each node and run the Local Clustering Coefficient (LCC) algorithm, which is implemented in [54].
Results (Figure 16): We find that, on the seven datasets (six for WBI), the running time of CuckooGraph on LCC is 2.06×2.06$\times$2.06 ×, 5.80×5.80$\times$5.80 ×, 3.94×3.94$\times$3.94 ×, and 4.21×4.21$\times$4.21 × faster than that of LiveGraph, Spruce, Sortledton, and WBI on average, respectively.
Analysis: Similar to the analysis in §$\S$ V-E4.